### v1

# Optimization under rare events: scaling laws for linear chance-constrained programs

Abstract.
We consider a class of chance-constrained programs in which profit needs to be maximized while enforcing that a given adverse event remains rare. Using techniques from large deviations and extreme-value theory, we show how the optimal value scales as the prescribed bound on the violation probability becomes small and how convex programs emerge in the limit. We also use our results to show that the popular CVaR approximation is asymptotically optimal under light-tail assumptions, but it is sub-optimal in a heavy-tailed setting. Then, we apply point process techniques and random set theory to study the suboptimality gap in the Monte Carlo based scenario approach.
Keywords: chance-constrained optimization, rare events, extreme values, regular variation.
MSC Class: 90C15 (primary), 60F10, 60G70 (secondary).
1. Introduction
We analyze a class of optimization problems where the goal is to maximize profit, subject to a reliability constraint that requires the probability of an undesirable event below a small level $\delta$. These formulations, known as chance-constrained (CC) optimization problems, arise in a wide range of applications, including:
  • Financial planning. Banks and insurance companies need to optimize their resources such that the probability of insolvency is kept small. For example, in fixed income portfolio optimization, an investment grade portfolio has a default rate of about $\delta=10^{-4}$, see Frank (2008).
  • Network design. In ultra-reliable communication system design for 5G wireless networks, the failure probability should be no bigger than $\delta=10^{-5}$, as is explained in Alsenwi et al. (2019).
  • Power grid operations. The European Union has set a goal that the amount of reserves available to mitigate possible frequency disturbance should be sufficient, up to once every 20 years. As these reserves are typically procured on an hourly basis, this amounts to a safety constraint of $\delta$ smaller than $10^{-5}$, see Mosca (2024).
Additional examples originate in the design of bridges and control of aircraft, cf. Shi and Holzapfel (2021), where the probability of a rare event is sometimes required to be below $\delta=10^{-9}$. For more examples, see Subramanyam (2020), Tong et al. (2022b).
Because of their direct interpretation and versatility to be applied to various important areas, chance-constrained optimization has been a subject of interest in Operations Research for several decades. It has become one of the standard formulations in risk-constrained optimization and is covered as such in several textbooks cf. Prekopa (1970), Prékopa (2003), Klein Haneveld et al. (2020), Shapiro et al. (2021). Unfortunately, chance-constrained problems are typically not very tractable: the set of feasible solutions may be non-convex, and evaluating the chance constraint can be a computational challenge. In general, CC problems are provably NP-hard, see Luedtke et al. (2010). As a result, researchers often resort to various approaches that provide bounds or relaxations to approximate the solution to CC optimization problems.
Our goal in this paper is to leverage the fact that in many applications, such as those described earlier, $\delta$ is very close to zero. We use this feature, together with large-deviations principles and extreme-value theory to gain insights into the asymptotic behavior (as $\delta\downarrow 0$) of a stylized family of CC problems and their relaxations. In order to provide a more detailed description of our results and their implications, let us provide a brief overview of various methods in CC optimization.
One method, introduced in Nemirovski and Shapiro (2006a), constructs a convex restriction of the feasible region by upper bounding the reliability constraint by the so-called conditional value at risk (abbreviated with CVaR; henceforth we refer to this as the CVaR approach). Another option, popular because of its ease of implementation, is the so-called scenario approach, introduced by Calafiore and Campi (2006). The scenario approach consists of replacing the probability constraint by several Monte Carlo-based scenarios, all of which need to satisfy the reliability constraint. If the undesirable event needs to be kept rare, the number of required samples is large, which can be prohibitive, though it is sometimes possible to reduce the number of samples using importance sampling techniques, see Nemirovski and Shapiro (2006b), Barrera et al. (2016), Blanchet et al. (2024).
Recently, several authors (see e.g. Subramanyam (2020), Nesti et al. (2019), Tong et al. (2022a)) have suggested a different method, namely to approximate the probability constraint using ideas from large-deviations theory. The approach can be described as follows. Suppose that ${\bf P}(G(x))\leq\delta$ is a safety constraint, $x$ is the optimization decision variable, and $G(x)$ is an undesirable event. A (heuristic) large-deviations approximation of ${\bf P}(G(x))$ can be obtained by replacing the integral of the density of an underlying probabilistic model (say $f$) over the set $G(x)$ by a supremum, i.e.:
${\bf P}(G(x))=\int_{G(x)}f(u)\mathrm{d}u\approx\sup_{u\in G(x)}f(u)=\exp\{-\inf_{u\in G(x)}\log f(u)\}.$ (1)
The previous approximation can often be made rigorous by introducing a suitable asymptotic regime. Using (1), the chance-constrained program can be approximated by a bi-level (or nested) optimization problem (involving $x$ at the top level and $u$ at the bottom level).
The papers Subramanyam (2020), Nesti et al. (2019), and Tong et al. (2022a) focus on various computational aspects of this bi-level procedure. The analysis in those papers requires sufficient tractability for the underlying density $f$, and this is carried out in situations where (for example) the randomness is a mixture of Gaussian distributions or where the randomness arises from another specific light-tailed distribution. Further, this analysis focuses on developing approximations as in (1) that result in computationally tractable procedures.
Moving on to a more precise description of our contributions, this paper takes a different perspective: we derive rigorous approximations like (1) allowing both light and heavy-tailed distributions; in the latter case, the analogues of (1) can be harder to formulate. In turn, we consider a sequence of CC problems indexed by the safety parameter $\delta$, and complete the following technical tasks:
  1. (a)
    we show how large-deviations approximations for the solution of CC problems emerge in the limit and analyze the scaling of the chance-constrained program’s optimal value $v_{\delta}$ as $\delta\downarrow 0$;
  2. (b)
    we compare $v_{\delta}$ both to the performance of the CVaR relaxation and the scenario approach;
  3. (c)
    we identify conditions for which the latter approaches are asymptotically optimal or not.
An attractive feature of our analysis is that it is rooted in multivariate extreme-value theory, allowing us to consider classes of distributions that are either light-tailed or heavy-tailed with a rich dependence structure. We restrict our analysis to a class of linear programs to keep it crisp. Specifically, the insights that we obtained based on our technical development can be summarized as follows:
  1. (1)
    If the underlying random variables are light-tailed, then we show (Theorem 1 below) that the basic formulation of the chance-constrained program, given in ($\rm{CCP}_{\delta}$) below, converges to a convex optimization problem as $\delta\downarrow 0$, and that an appropriate version of (1) emerges. While this leads to a characterization of the growth rate of $v_{\delta}$, it is nontrivial to construct a sequence of asymptotically optimal solutions from the limiting optimization problem: essentially, the limit of the logarithmic estimate (1) is rather rough to guarantee feasibility for fixed $\delta>0$. However, the asymptotics are sharp enough to derive solutions with any prescribed (non-negative) asymptotic optimality gap.
  2. (2)
    Also in the context of light-tailed distributions, we show that both the CC problem and its CVaR relaxation converge to the same convex program in the regime $\delta\downarrow 0$. Thus, CVaR provides a sequence of convex approximations that are both conservative and asymptotically optimal: their optimal values scale in the same way as the actual profit. For a precise formulation, see Theorem 2. In short, for light-tailed settings, the CVaR relaxation is asymptotically exact.
  3. (3)
    A similar, but more nuanced conclusion, once again in the light-tailed environment, holds for the scenario approach: in this case, we show that the sequence of random linear programs, indexed by the number of scenarios, converges, after appropriate normalization, to a deterministic optimization problem as the number of scenarios is large. However, it must be emphasized that to match the behavior of ($\rm{CCP}_{\delta}$), the scenario approach requires a careful choice in the number of scenarios $k$ as function of $\delta$, which differs from the standard prescription provided by Calafiore and Campi (2006); see also (6) below. Again, feasibility for the original problem cannot be guaranteed in general. If we take the standard recommendation for $k$ as given in Calafiore and Campi (2006) to guarantee feasibility with high probability, the gap between the actual optimal profit and the optimal profit resulting from the scenario approach diverges. We refer to Theorem 3 and its discussion for details.
  4. (4)
    If the underlying random variables are heavy-tailed (more precisely, if their tails are of regular variation), we show (Theorem 4) that our basic chance-constrained problem converges to a convex optimization problem after appropriate renormalization. It turns out that (1) is not adequate here: essentially logarithmic asymptotics as in (1) are not enough; a more precise asymptotic estimate is needed. Fortunately, such an estimate is available in the heavy-tailed case, using multivariate regular variation.
  5. (5)
    A normalized sequence of CVaR constraints converges to a similar limit as the chance constraints, except with a different value in a certain constant. This makes CVaR asymptotically suboptimal in the heavy-tailed case. For example, if the underlying random variables have a power law with index $\alpha$, CVaR yields a value of the objective function which is a factor $1-1/\alpha$ of the actual optimum, as $\delta\downarrow 0$. A detailed statement is Theorem 5 below.
  6. (6)
    The scenario approach behaves even worse in the heavy-tailed case: the appropriately normalized sequence of random optimization problems converges to a limit, but this limit is still random, cf. Theorem 6 below. This is in contrast to methods such as sample-average approximation (also known) as empirical optimization in which, as the Monte Carlo sample size increases, the sample-based approximation converges to the desired solution. Our results suggest that in the heavy-tailed case the scenario approach must be modified to obtain asymptotically optimal solutions. We utilize concepts from point-process theory by Resnick (2007) and random set theory, see Molchanov (2017), to establish these results.
The rest of this paper is organized as follows. We define our model and state our assumptions in Section 2. In Section 3, we derive our results for light-tailed distributions. Our results for heavy-tailed distributions can be found in Section 4. In Section 5 we consider an example showing how it is possible to model dependence between random variables in our framework for light-tailed distributions, and how using approximations like (1) impacts feasibility. Proofs of our results can be found in Section 7, which also develops a general-purpose result that circumvents the use of epi-convergence.
2. Problem description
Basic formulation. Consider the problem
$\mbox{maximize }c^{T}x\hskip 28.45274pt\mbox{ subject to }x\in X,\mbox{ and }{\bf P}(\phi(x,L)>1)\leq\delta,$ ($\rm{CCP}_{\delta}$)
where $x\in\mathbb{R}^{m}$ is an $m$-dimensional decision vector, $c$ is a vector with non-negative entries, $X$ is a compact convex subset of $[0,\infty)^{m}$ which contains the set $[0,h]^{m}$ for some $h>0$, and $L\geq 0$ is an $n$-dimensional random vector. The elements of $L$ are often referred to as risk factors and may be dependent. We assume there exist $n\times m$ matrices $A^{(i)},i=1,...,d$ with non-negative entries and set
$\phi(x,L):=\sup_{i=1,...,d}x^{T}A^{(i)}L.$ (2)
This formulation is general enough to model linear programs with random coefficients in the constraints.
As mentioned earlier, the parameter $\delta>0$ denotes the risk tolerance. We focus on the case where the reliability constraint is a rare event, i.e. $\delta\downarrow 0$. We aim to understand how the optimal profit scales as $\delta\downarrow 0$.
Definition 1 (Asymptotic optimality).
[…]
$\limsup_{\delta\downarrow 0}{\bf P}(\phi(x_{\delta},L)>1)/\delta\leq 1.$ (3)
We call $(x_{\delta})_{\delta>0}$ asymptotically optimal if
$c^{T}x_{\delta}\,\big{/}\,v_{\delta}\rightarrow 1,\qquad\text{as }\delta\downarrow 0.$ (4)
We now introduce two formulations that are popular alternatives to ($\rm{CCP}_{\delta}$); we are interested in whether their solutions are asymptotically optimal. The determination of asymptotic optimality is facilitated by Theorems 1 and 4 below, as they establish the asymptotic scaling of ($\rm{CCP}_{\delta}$) for light-tail and heavy-tail distributions, respectively.
CVaR formulation. The problem ($\rm{CCP}_{\delta}$) is generally not convex. For this reason, Nemirovski and Shapiro (2006a) introduce a convex relaxation which typically (in particular in the case of non-atomic distributions) coincides with the conditional value-at-risk
${\bf CVaR}(\phi(x,L))=\inf_{\tau\in{\mathbb{R}}}\Big{[}\tau+\frac{1}{\delta}{\bf E}[(\phi(x,L)-1-\tau)^{+}]\Big{]}.$ (5)
In particular, they propose to solve the optimization problem
$\mbox{maximize }c^{T}x\hskip 28.45274pt\mbox{ subject to }x\in X,\mbox{ and }{\bf CVaR}(\phi(x,L))\leq 0.$ ($\rm{CVaR}_{\delta}$)
Let $v_{{\bf CVaR},\delta}$ be the optimal solution of the CVaR version of ($\rm{CCP}_{\delta}$). The CVaR constraint in ($\rm{CVaR}_{\delta}$) is more conservative than the chance constraint in ($\rm{CCP}_{\delta}$), so $v_{{\bf CVaR},\delta}\leq v_{\delta}$. However, as this optimization problem is convex, it may be more tractable than ($\rm{CCP}_{\delta}$). Theorems 2 and 5 below show how $v_{{\bf CVaR},\delta}$ compares to $v_{\delta}$ as $\delta\downarrow 0$.
Scenario formulation. Another approach towards solving chance-constrained programs is to resort to Monte-Carlo methods, where the chance constraint is replaced by constraints of the form $\phi(x,L^{({i})})\leq 1$, with $L^{(1)},...,L^{(k)}$ a number of samples called scenarios. This results in the optimization problem
$\mbox{maximize }c^{T}x\hskip 28.45274pt\mbox{subject to }x\in X\mbox{ and }\phi(x,L^{(i)})\leq 1,\quad\forall i\in[k],$ ($\mathrm{SCP}_{k}$)
where $[k]:=\{1,\ldots,k\}$. To ensure feasibility of ($\rm{CCP}_{\delta}$) with high enough probability, $k$ needs to be chosen appropriately as $\delta\downarrow 0$. In Calafiore and Campi (2006) it is shown that the choice
$k(\delta)=\frac{2}{\delta}\log\frac{1}{\beta}+2d+\frac{2d}{\delta}\log\frac{2}{\delta}$ (6)
gives a solution that is feasible with probability at least $1-\beta$. This bound is valid for any function $\phi$. We show in Theorems 3 and 6 below that it leads to conservative solutions in our setting: letting $X^{k(\delta)}$ be a solution of ($\mathrm{SCP}_{k}$) with $k=k(\delta)$, then the random variable $c^{T}X^{k(\delta)}/v_{\delta}$ always converges to 0 in probability under our assumptions.
Plan. We analyze the above three families of optimization problems in the next two sections by making appropriate assumptions on $L$. We will apply concepts and techniques from multivariate extreme-value theory and large-deviations theory. We begin with the case in which the elements of $L$ are light-tailed.
3. Light tails
In this section, we analyze ($\rm{CCP}_{\delta}$), ($\rm{CVaR}_{\delta}$), and ($\mathrm{SCP}_{k}$) for a large class of distributions guaranteeing that all moments of $L$ are finite. In particular, we assume that $Q(x):=-\log{\bf P}(L>x)$ is multivariate regularly varying: there exists a constant $\beta>0$, a slowly varying function $h$ and a non-negative continuous function $\lambda:[0,\infty)^{n}\mapsto\mathbb{R}$ such that for $x\in[0,\infty)^{n}$, with $q(r)=h(r)r^{\beta}$,
$Q(rx)/q(r)\rightarrow\lambda(x),\quad\text{as }r\rightarrow\infty.$ (7)
The convergence as stated above is pointwise. However, it turns out to hold uniformly on compact sets because of regular variation, see Resnick (2007). Note that if we divide by $|x|$ and use that $h(r|x|)/h(r)\rightarrow 1$ as $r\rightarrow\infty$ for $|x|>0$, we obtain
$\lambda(x)=\lambda(x/|x|)\times|x|^{\beta}.$ (8)
This family of distributions includes multi-variate normal distributions (for which $\beta=2$), exponential and gamma distributions ($\beta=1$) and Weibull distributions with a sub-exponential tail ($\beta<1$). The dependence structure can be rather arbitrary: the main form of dependence which is relevant is to which extent large values in the sample are correlated, which is encoded in $\lambda$. We give an example of $\lambda$ in Section 5.
We now apply large-deviations theory to approximate ($\rm{CCP}_{\delta}$) for small $\delta$. Set for $b\in[0,\infty)^{n}$
$I(b):=\inf_{x}\big{\{}\lambda(x):bx\geq 1\big{\}}=|b|^{-\beta}I(b/|b|).$ (9)
In Section 7.1 we show that the set $\{b:I(b)\geq 1\}$ is convex, compact, possesses a non-empty interior, and
$-\log{\bf P}(b^{T}L>r)\sim I(b)q(r).$ (10)
Here, $\sim$ means that the ratio of both functions converges to $1$ as $r\rightarrow\infty$. Next, we use the principle of the largest term (see e.g. Ganesh et al. (2004), Lemma 2.1) to approximate our chance constraint by
$-\log{\bf P}(\max_{i}y^{T}A^{(i)}L>r)\sim\min_{i}I(y^{T}A^{(i)})q(r)=:J(y)q(r),$ (11)
making the change of variable $x=y/r$. Now, consider the problem
maximize $\mbox{maximize }c^{T}y\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }J(y)\geq 1.$ ($\rm{LT}$)
Formulation ($\rm{LT}$) is motivated by making the change of variable $x=y/r$ in ($\rm{CCP}_{\delta}$), replacing the probability in ($\rm{CCP}_{\delta}$) by the approximation ${\bf P}(\max_{i}y^{T}A^{(i)}L>r)=\exp(-J(y)q(r)(1+o(1))$, and choosing $r=q^{-1}(\log 1/\delta)$. This ensures that the chance constraint, after taking logarithms, has the desired behavior. To make this work, we assume that $q$ is continuous for large $r$.
The problem ($\rm{LT}$) is tractable: using Sion’s minimax theorem, cf. Sion (1958), strong duality holds, and since $\{J(y)\geq 1\}$ has non-empty interior, ($\rm{LT}$) can be solved by standard Lagrangian methods. In Section 5, we give examples for $\lambda$ for which $I$ itself has an explicit expression.
Our first result is that ($\rm{LT}$) determines the correct growth rate of $v_{\delta}$:
Theorem 1.
[…]
To prove this result, a verification of the pointwise convergence property (11) is not sufficient. The standard way in the optimization literature to prove a result of the type of Theorem 1 is through epi-convergence, see Rockafellar and Wets (1998). In our proof, we circumvent this and follow a more direct approach; we refer to Section 7 for details.
Based on Theorem 1, it is tempting to conjecture that $y^{*}/q^{-1}(\log 1/\delta)$ is asymptotically optimal, but this may not be the case, as the asymptotic estimate (11) is not sharp enough to guarantee the asymptotic feasibility condition (3), as is exemplified in Section 5. One may instead consider $(1-\eta)y^{*}/q^{-1}(\log 1/\delta)$, which is asymptotically feasible for any $\eta>0$ (this follows from the proof of the next theorem in Section 7.2), and accept an optimality gap $\eta$.
In contrast, ($\rm{CVaR}_{\delta}$) produces an asymptotically optimal sequence of solutions:
Theorem 2.
$v_{{\bf CVaR},\delta}/v_{\delta}\rightarrow 1$ as $\delta\downarrow 0$.
The reason why this is true is that the overshoot $\phi(y^{*}/q^{-1}(\log 1/\delta),L)-1$, conditional on $\phi(y^{*}/q^{-1}(\log 1/\delta)/,L))>1$, converges to 00. In addition, the optimizing value of $\tau$ in (5) converges to 00 as $\delta\downarrow 0$. Together, these two properties imply that the constraint in the CVaR formulation ($\rm{CVaR}_{\delta}$) behaves asymptotically the same as the probability in ($\rm{CCP}_{\delta}$). This makes the CVaR formulation appealing when optimizing under rare event constraints in a light-tail setting.
To assess ($\mathrm{SCP}_{k}$) in the region $k\rightarrow\infty$, let $L^{(j)},j=1,...,k$ be $k$ i.i.d. samples of $L$. Using the insights from Theorem 1, this motivates us to consider a scaled problem. We make a similar change of variables $x=y/q^{-1}(\log k)$. We exploit that $\phi(y/r,L)=\phi(y,L)/r$ and consider
$\mbox{maximize }c^{T}y\hskip 14.22636pt\mbox{subject to }\phi(y,L^{(i)})\leq q^{-1}(\log k),\quad\forall i\in[k].$ (${\rm{SCP}\textnormal{-}\rm{LT}}_{k}$)
Theorem 3.
[…]
To recover the optimal growth rate for the optimal value $v_{\delta}$ of ($\rm{CCP}_{\delta}$), we see from Theorem 1 that it is required that $k$ is proportional to $1/\delta$. As the choice $k(\delta)$ to guarantee feasibility diverges faster than $1/\delta$, the choice $k=1/\delta$ may lead to an asymptotically infeasible solution.
Summarizing our findings in the light-tail case, we see that the behavior of the optimal value $v_{\delta}$ can be characterized using the rare-event approximation ($\rm{LT}$). Because this approximation is based on logarithmic asymptotics (11), the rescaled optimum of $y^{\ast}$ may not be feasible, even for small $\delta$.
The sampling approach leads to a random solution $Y^{(k)}$, which, after appropriate normalization, converges to the solution of ($\rm{LT}$). Apart from the computational challenges that play a role if $k$ is large, this procedure leads to a tension between feasibility and optimality: to get a feasible solution, the choice $k(\delta)$ leads to an optimal profit which is of smaller order than $v_{\delta}$ as $\delta\downarrow 0$.
Finally, the CVaR approach, while not as tractable as ($\rm{LT}$), and more conservative than ($\rm{CCP}_{\delta}$), is asymptotically optimal in a light-tail regime, reinforcing the fact that CVaR is one of the main approaches for solving chance constraint programs. In the next section, we will examine if these findings still hold if the underlying random variables are heavy-tailed.
4. Heavy tails
Throughout this section, we suppose that $L$ has a distribution that is of multivariate regular variation, which covers the case in which the elements of $L$ have a power law. To describe our assumptions in a convenient form, we use a representation which (apart from not taking logarithms) is different from the one in Section 3, building on an equivalent characterization of regular variation developed in Basrak et al. (2002)). Specifically, let ||$\cdot$|| ⋅ | be the ${\mathcal{L}}_{1}$ norm. Then $\bar{F}(r)={\bf P}(|L|>r)=h(r)r^{-\alpha},\alpha>1$, with $h$ slowly varying (i.e. $h(ar)/h(r)\rightarrow 1$ as $r\rightarrow\infty$ for $a>0$) and
${\bf P}\big{(}L/|L|\in\cdot\,\big{|}\,|L|>r\big{)}\rightarrow{\bf P}(\Theta\in\cdot),\hskip 28.45274ptr\rightarrow\infty.$ (12)
The idea behind this representation is that, if we transform $L$ in polar coordinates, the radius and angle become independent, conditional upon the radius getting large. This formulation allows dependence between extreme values of risk factors. For background on multivariate regular variation, see Basrak et al. (2002), Resnick (2007).
We now construct an appropriately normalized version of ($\rm{CCP}_{\delta}$) in the regime $\delta\downarrow 0$. Set $x=y/r_{\delta}$, with $r_{\delta}:=\bar{F}^{-1}(\delta)=\inf\{r:{\bf P}(|L|>r)\leq\delta\}$. Now, write the constraint in ($\rm{CCP}_{\delta}$) as ${\bf P}(\phi(y,L)>r_{\delta})\leq\delta.$ We assume that $\bar{F}(t)$ is eventually strictly monotone and continuous which makes the constraint equivalent to
${\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})\leq 1.$ (13)
Letting $\delta\downarrow 0$, using (12), we show in Section 7.4 that ${\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})\rightarrow{\bf E}[\phi(y,\Theta)^{\alpha}]$, and that ($\rm{CCP}_{\delta}$) can be approximated by
maximize $\mbox{maximize }y^{T}c\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }{\bf E}[\phi(y,\Theta)^{\alpha}]\leq 1.$ ($\rm{HT}$)
An appealing feature of ($\rm{HT}$) is that the feasible region of this problem is convex, which generally is not the case for ($\rm{CCP}_{\delta}$). We prove the following theorem in Section 7.4:
Theorem 4.
[…]
Unlike in the light-tail case, the value $y^{*}$ leads to an asymptotically optimal solution as asymptotic feasibility is guaranteed. In contrast, in the heavy-tailed case ($\rm{CVaR}_{\delta}$) is no longer asymptotically optimal:
Theorem 5.
$v_{{\bf CVaR},\delta}/v_{\delta}\rightarrow 1-1/\alpha$ as $\delta\downarrow 0$.
The reason behind the subobtimality is that the overshoot $\phi(y,L)-r_{\delta}$, conditioned on $\phi(y,L)>r_{\delta}$ is not of smaller order than $r_{\delta}$ when $\phi(y,L)$ has a heavy tail.
In the previous section, we observed that the sample approximation ($\mathrm{SCP}_{k}$) of ($\rm{CCP}_{\delta}$) has some optimality properties in the light-tail case, since the space of feasible solutions converges, after appropriate normalization, to a deterministic set which coincides with the feasible region of the limiting optimization problem ($\rm{LT}$). In the heavy-tail case, this is no longer true. In contrast, a random solution emerges, making any optimality property elusive.
To see this, consider the simplest possible case, where the goal is to maximize a scalar $x$ subject to the constraint that ${\bf}P(xL>1)\leq\delta$, with $L$ also scalar. ($\mathrm{SCP}_{k}$) now can be formulated as maximizing $x$ subject to $x\max_{i=1}^{k}L^{(i)}\leq 1$. If $L$ is light-tailed, we can normalize $\max_{i=1}^{k}L^{(i)}$ such that it converges to a deterministic constant (we have done this in a more general setting to obtain Theorem 3). However, in the heavy-tail case, such a normalization is not possible. In fact, $\max_{i=1}^{k}L^{(i)}/\bar{F}^{-1}(1/k)$ converges in distribution to a random variable with a Fréchet distribution, see Resnick (2007), Nair et al. (2022) for background.
To describe our general technical result we use point-process machinery commonly used in extreme-value theory, cf. Resnick (2007) as well as random set theory as developed in (Molchanov and Strokorb (2016)). Let $N$ be a Poisson random measure on the space $(0,\infty]\times S_{n-1}^{+}$, where $S_{n-1}^{+}$ is the intersection of the unit sphere and the positive orthant.
Let $N$ be a Poisson random measure on $(0,\infty]\times S_{n-1}^{+}$ with intensity measure $\zeta(\cdot)=\nu_{\alpha}(\cdot)\times{\bf P}(\Theta\in\cdot)$, where $\nu_{\alpha}$ has density $\alpha u^{-\alpha-1}$. This should be interpreted as follows: if $F$ is a subset of $(0,\infty]\times S_{n-1}^{+}$, then $N(F)$ has a Poisson distribution with parameter $\zeta(F)$.
Under our assumptions, the point process (with $\epsilon$ being the Dirac measure)
$N_{k}(\cdot)=\sum_{i=1}^{k}\epsilon_{|L^{(i)}|/\bar{F}^{-1}(1/k),L^{(i)}/|L^{(i)}|}(\cdot)$ (14)
is converging in distribution to $N$; this is a direct application of Theorem 6.2 of Resnick (2007). Similar to (${\rm{SCP}\textnormal{-}\rm{LT}}_{k}$) we perform a change of variables $x=y/\bar{F}^{-1}(1/k)$, exploit the homogeneity of $\phi$ and define the random variable $V_{SCP}(k)$ to be the optimal value of the optimization problem
$\mbox{maximize }c^{T}y\hskip 28.45274pt\mbox{subject to }y\geq 0\mbox{ and }\phi(y,L^{(i)})\leq\bar{F}^{-1}(1/k),\quad\forall i\in[k].$ (${\rm{SCP}\textnormal{-}\rm{HT}}_{k}$)
In Section 7.6 we apply concepts from random set theory (cf. Molchanov (2017)) to show
Theorem 6.
$V_{SCP}(k)$ converges weakly to the optimal value $V^{*}$ of the optimization problem
$\mbox{maximize }c^{T}y\hskip 28.45274pt\mbox{subject to }y\geq 0\mbox{ and }\phi(y,z)\leq 1,\quad\forall z\in N.$ (15)
In conclusion, comparing the three main methods to (approximately) solve ($\rm{CCP}_{\delta}$) in the heavy-tailed case, the large-deviations approximation ($\rm{HT}$) yields an asymptotically optimal solution, and is characterized by a convex optimization problem. In contrast, the CVaR formulation yields a solution that stays a factor $1-1/\alpha$ away from the optimum, while the scenario approach does not converge to a deterministic solution or approximation of ($\rm{CCP}_{\delta}$) at all.
5. Example for light-tailed distributions
While the heavy-tail case directly leads to the convex program ($\rm{HT}$), the light-tail case leads to a bi-level optimization problem involving the function $I$ in (9), making it less straightforward to parse. In this section, we provide a computationally tractable example for the light-tail case that also provides insights into how the tail behavior of $L$ and its dependence structure may impact the optimal solution of ($\rm{CCP}_{\delta}$) as $\delta\downarrow 0$. Throughout this section, we assume $d=1,m=n$ and consider the problem
$\mbox{maximize }c^{T}x\hskip 28.45274pt\mbox{ subject to }x\geq 0,\mbox{ and }{\bf P}\Big{(}\sum_{i=1}^{n}x_{i}a_{i}L_{i}\geq 1\Big{)}\leq\delta,$ (16)
for some constants $a_{i}>0,i=1,...,n$. Suppose that $L$ satisfies (7), with $\lambda$ given by
$\lambda(x):=\ell(x^{\beta}):=\ell(x_{1}^{\beta},...,x_{n}^{\beta}):=\bigg{(}\sum_{i=1}^{n}x_{i}^{\beta\theta}\bigg{)}^{1/\theta},\hskip 28.45274pt\theta\geq 1.$ (17)
Here, $\ell(x)$ corresponds to the so-called Gumbel-Hougaard copula (see e.g. Segers (2012)). For $\theta=1$, large coordinates in $L$ occur independently; this is no longer true if $\theta>1$. For $\theta=\infty$, we take $\ell(x)=\max_{i}x_{i}$, so that large values are perfectly correlated.
If $\gamma=\beta\theta\leq 1$, then $I(b)=\min\{(\sum_{i=1}^{n}x_{i}^{\gamma})^{1/\theta}:b^{T}x\geq 1\}$ has as optimal solution $x_{i^{*}}=1/b_{i^{*}}$ with $i^{*}=\arg\max b_{i}$ (with arbitrary tie-breaking rule) and $x_{i}=0$ for $i\neq i^{*}$. Consequently,
$I(b)=|b|_{\infty}^{-\beta}.$ (18)
Thus, if $\gamma\leq 1$, the limiting approximation ($\rm{LT}$) of ($\rm{CCP}_{\delta}$) simplifies to
$\mbox{maximize }c^{T}y\hskip 28.45274pt\mbox{ subject to }y\geq 0,\mbox{ and }\max_{i}a_{i}y_{i}\leq 1,$ (19)
which has solution $y_{i}^{\ast}=1/a_{i}$.
If $\gamma=\beta\theta>1$, then we find using the Karush-Kuhn-Tucker (KKT) conditions that the optimal $x_{i}$ in the definition of $I(b)$ equals $b_{i}^{1/(\gamma-1)}/\sum_{j}b_{j}^{\gamma/(\gamma-1)}$, and after some simplifications we obtain
$I(b)=\left(\sum_{i=1}^{n}b_{i}^{\gamma/(\gamma-1)}\right)^{-(\gamma-1)/\theta}.$ (20)
Thus, if $\gamma>1$, the limiting approximation ($\rm{LT}$) of ($\rm{CCP}_{\delta}$) simplifies to
$\mbox{maximize }c^{T}y\hskip 28.45274pt\mbox{ subject to }y\geq 0,\mbox{ and }\sum_{i=1}^{n}(a_{i}y_{i})^{\gamma/(\gamma-1)}\leq 1.$ (21)
Using the KKT conditions we obtain the solution
$y_{i}^{\ast}=(c_{i}/a_{i})^{\gamma-1}/\Big{(}\sum_{j=1}^{n}(c_{j}/a_{j})^{\gamma}\Big{)}.$ (22)
If the coordinates of $L$ are independent in the tail ($\theta=1)$, then the nature of $I(b)$ changes at the point $\beta=1$ which is precisely the point where the moment-generating function of $L$ becomes finite. For $\beta<1$, the tails of $L_{i}$ are heavy-tailed Weibull, for which it is known that large values of sums are caused by a single big jump, leading to the simple form of the optimal solution.
Interestingly, the impact of the dependence structure, summarized by $\theta$, is nontrivial, in the sense that the single-big jump intuition is not relevant in our decision-making context if the tail of $L_{i}$ is not very heavy, and the dependence is sufficiently strong, i.e. if $\beta>1/\theta$.
We mentioned before that asymptotic optimality in the light-tailed case is not guaranteed as the solution of ($\rm{LT}$) after normalization may not be feasible for the problem ($\rm{CCP}_{\delta}$). To illustrate this in more detail in the context of our example, consider the case where $\theta=1$ and $\beta<1$. Take also $q(u)=u^{\beta}$. In this case, the chance constraint in (16), taking the scaled solution $x_{i}=1/(a_{i}q^{-1}(\log 1/\delta))$, leads to the probability ${\bf P}\big{(}\sum_{i=1}^{n}L_{i}>q^{-1}(\log 1/\delta)\big{)}$. The distribution of $L$ is sub-exponential and we have the estimate, as $\delta\downarrow 0$,
${\bf P}\Big{(}\sum_{i=1}^{n}L_{i}>q^{-1}(\log 1/\delta)\Big{)}\sim n{\bf P}\big{(}L_{1}>q^{-1}(\log 1/\delta)\big{)}=n\delta,$ (23)
cf. Chapter 3 in Nair et al. (2022). Thus, our choice leads to a violation probability of a factor $n$ larger than it should be to satisfy (16).
For $\beta=1$ we obtain a sum of i.i.d. exponential random variables, for which
${\bf P}\Big{(}\sum_{i\in[n]}L_{i}>q^{-1}(\log 1/\delta)\Big{)}$ $\sim{\bf P}\big{(}\mathrm{Gamma}(n,1)>(\log 1/\delta)\big{)}$ (24)
$=\int_{u=(\log 1/\delta)}^{\infty}\frac{1}{\Gamma(n)}x^{n-1}{\mathrm{e}}^{-x}\mathrm{d}u$ (25)
$=\frac{1}{n!}\mathrm{e}^{-(\log 1/\delta)}(\log 1/\delta)^{n}=\delta\frac{(\log 1/\delta)^{n}}{n!}.$ (26)
This behavior is even worse compared to $\beta<1$ as $\delta\downarrow 0$.
6. Concluding comments
This paper provides the first systematic large-deviations analysis of chance-constrained optimization problems. To facilitate the development of our algorithmic and structural insights, we focus on a linear objective and a safety constraint controlling (i.e. upper bounding by a fixed constant) the $1-\delta$ quantile of the maxima of bi-linear functions (bi-linear both in the multivariate risk factor and the decision parameter). We distinguish two regimes according to the tails of the risk factor, namely, light and heavy tailed regimes.
Under the light-tail assumption, the main takeaway from Section 3 is that the CVaR formulation is a good approach to chance-constrained optimization in a rare-event regime. The scenario approach and the large-deviations approximation (both suitably modified as explained in Section 3) can be a good approach to CC optimization, though for those methods, there is a nontrivial trade-off between feasibility and optimality.
In the heavy-tailed regime, we explicitly show that the CVaR relaxation is no longer asymptotically optimal: its solution remains a factor $1-1/\alpha$ away from the optimum. The scenario approach behaves even worse, as explained in Section 4. In contrast, the large-deviations approximation is asymptotically optimal in this case, as it is based on a sharper version of the large-deviations approximation (1).
There are several natural research questions that may be considered: (i) we expect that large-deviations approximations, for both light-tails and heavy-tails may still yield convex programming problems for functions $\phi$ which may not be bilinear (ii) computational considerations, it will be of interest to consider scaling regimes where also the dimension of the decision variable $x$ and/or the dimension of $L$ may grow; we kept it fixed in this study; (iii) the large-deviations insights we developed could be used to develop new algorithms for chance-constrained optimization which are asymptotically optimal in the rare event regime, and also produce close to optimal solutions for fixed values of $\delta$, just like the notion of asymptotic efficiency has led to novel algorithms in the area of rare-event simulation.
7. Proofs
We prove our theorems using the following generic convergence result. There are related results in the literature, cf.  Bonnans and Shapiro (2000), Rockafellar and Wets (1998), but our direct approach seems the most efficient for our purposes.
Proposition 1.
Assume that $c\geq 0$, $g_{\ell}$ is non-decreasing in $y$ for each $\ell$ and that $g_{\ell}\rightarrow g$ pointwise as $\ell\rightarrow\infty$, with $g$ continuous and strictly increasing. Let $v_{*}=\sup\{c^{T}y:g(y)\leq 1,y\geq 0\}$. Further, define $v_{\ell}=\sup\{c^{T}y:g_{\ell}(y)\leq 1,y\geq 0$}. Then, $v_{\ell}\rightarrow v_{*}$.
If $c=0$ the result is immediate. So, we assume that at least one entry of $c$ is positive. We first consider the case $v_{*}=\infty$. For every $a>0$ there exists $y\geq 0$ such that $c^{T}x>a$. By continuity of $g$ there exists $\gamma\in(0,1)$ such that $g(\gamma y)<1$ and $c^{T}y>a/2$. This implies that $g_{\ell}(\gamma y)\leq 1$ and therefore $v_{\ell}\geq a/2$, for all $\ell$ large enough. Thus, we may assume that $v_{*}<\infty$. Next, observe that for each $\eta<1$, $g_{\ell}(\eta y^{*})<1$ for $\ell$ sufficiently large. Consequently, $c^{T}y_{\ell}\geq\eta c^{T}y^{*}$ for $\ell$ sufficiently large, and by letting $\eta\uparrow 1$ we obtain that
$\liminf_{\ell\rightarrow\infty}c^{T}y_{\ell}\geq c^{T}y^{*}.$ (27)
To prove a corresponding asymptotic upper bound, we assume to the contrary that for some $\varepsilon>0$ and a sub-sequence $(v_{\ell})$ (which we also index by $\ell$) we have $v_{\ell}\geq v_{*}+2\varepsilon$.
We want to contradict the optimality of $y^{*}$. Let us partition $c=(\bar{c},0)$, where the part $\bar{c}$ has strictly positive elements. We then note that $v_{*}=\max\{\bar{c}^{T}\bar{y}:g(\bar{y},0)\leq 1,\bar{y}\geq 1\}$. Because $g_{\ell}$ is non-decreasing, we have that $v_{\ell}=\max\{\bar{c}^{T}\bar{y}:g_{\ell}(\bar{y},0)\leq 1,\bar{y}\geq 0\}.$ So, we may assume without loss of generality that $c$ has strictly positive elements. Now, let $y_{\ell}$ be an optimal solution corresponding to $v_{\ell}$ (if the solution is not achieved, we can work with an $\varepsilon/2$ solution). Since $g_{\ell}$ is non-decreasing, we can make the coordinates of $y_{\ell}$ smaller, obtaining $y^{\prime}_{\ell}$ which remains feasible and $c^{T}y^{\prime}_{\ell}=c^{T}y^{*}+\varepsilon$. Since the elements of $c$ can be assumed to be strictly positive, the set $C=\{x:c^{T}y=c^{T}y^{*}+\varepsilon\}$ is compact. Since the $g_{\ell}$’s are non-decreasing, the convergence $g_{\ell}\rightarrow g$ occurs uniformly on compact sets. By choosing $\ell$ sufficiently large we can construct $y^{\prime\prime}_{\ell}$ (possibly by reducing the coordinates of $y^{\prime}_{\ell}$) such that $g(y^{\prime\prime}_{\ell})\leq 1$, contradicting the optimality of $y^{*}$. This completes the proof of this proposition. $\square$
7.1. Proof of Theorem 1
We first show that the feasible region of ($\rm{LT}$) is convex and compact. For this, it suffices to show that $\{b:I(b)\geq 1\}$ is convex and compact, where $I$ is defined in (9). Let $b^{(1)}$ and $b^{(2)}$ be such that $I(b^{(i)})\geq 1,i=1,2$. Let $a\in(0,1)$. For the desired convexity property, we have to show that $I(ab^{(1)}+(1-a)b^{(2)})\geq 1$. Since $\lambda$, defined in (7), is continuous, non-decreasing, and satisfies the relation $\lambda(rx)=r^{\beta}\lambda(x)$, there exists an $x^{(a)}\geq 0$ such that $I(ab^{(1)}+(1-a)b^{(2)})=\lambda(x^{(a)})$. Since $(ab^{(1)}+(1-a)b^{(2)})x^{(a)}\geq 1$, we have that either $b^{(1)}x^{(a)}\geq 1$ or $b^{(2)}x^{(a)}\geq 1$. Suppose for concreteness that $b^{(1)}x^{(a)}\geq 1$. Then, $I(ab^{(1)}+(1-a)b^{(2)})\geq I(b^{(1)})\geq 1$ implies the desired convexity property. In fact, a similar argument as that provided in the proof of Proposition 1, shows that $I(\cdot)$ is continuous (using that the minimization defining $I(b)$ can be assumed to be taken over a compact set); this implies that $\{b:I(b)\geq 1\}$ is closed. The compactness of the set $\{b:I(b)\geq 1\}$ follows from the relation $\lambda(rx)=r^{\beta}\lambda(x)$, $\beta>0$.
Next, we show that (10) holds, using standard textbook arguments from large-deviation theory, see e.g. Dembo and Zeitouni (1998): note that, if $B_{\delta}(x)$ is the $L_{\infty}$ ball of radius $\delta$ around $x$, the assumption on $L$ implies that $-\log{\bf P}(L/r\in B_{\delta}(x))\sim\inf_{u\in B_{\delta}(x)}\lambda(u)$. Therefore, using also the continuity of $\lambda$, the same estimate holds for the closure of $B_{\delta}(x)$, from which it follows that $L/r$ satisfies a weak large-deviations principle with rate function $\lambda$ and speed $q(r)$. Since $\lambda(ax)=a^{\beta}\lambda(x)$, $\lambda$ has compact level sets. Therefore, $L/r$ satisfies a full large-deviations principle with rate function $\lambda$ and speed $q(r)$. Since $b^{T}L$ is a continuous function of $L$ and $\lambda$ is continuous, (10) follows from the contraction principle.
We now turn to the proof of Theorem 1. We apply Proposition 1 with $\ell=1/\delta$, $g_{\ell}(x)=(\log\delta)/\log{\bf P}(\phi(x,L)>q^{-1}(\log 1/\delta))$, and $g(x)=1/J(x)$. The pointwise convergence $g_{\ell}(x)\rightarrow g(x)$ follows from (10). The fact that $g$ is strictly increasing follows from the homogeneity property of $J$. $\square$
7.2. Proof of Theorem 2
Since $v_{{\bf CVaR},\delta}\leq v_{\delta}$, it suffices to prove an asymptotic lower bound. For this, we utilize Theorem 1. Let $y^{*}$ be a solution of ($\rm{LT}$) and for fixed $\gamma>0$, define $y:=(1-\gamma)y^{*}$. Moreover, let $r_{\rm{LT},\delta}:=q^{-1}(\log 1/\delta)$. We will prove that $x=y/r_{\rm{LT},\delta}$ will be feasible for ($\rm{CVaR}_{\delta}$), for $\delta$ sufficiently small.
The optimal value $\tau=\tau_{\delta}$ in (5) with $x=y/r_{\rm{LT},\delta}$ solves a first-order optimality condition given by
$\delta={\bf P}\big{(}\phi(y,L)>r_{\rm{LT},\delta}(\tau_{\delta}+1)\big{)}.$ (28)
Since this equality is true for all $\delta>0$, this implies that $r_{\rm{LT},\delta}(\tau_{\delta}+1)\to\infty$. Combining this with the tail behavior of $\phi(y,L)$ given in Equation (11), we see that as $\delta\downarrow 0$
$\log(1/\delta)=(1+o(1))J(y)q\big{(}r_{\rm{LT},\delta}(\tau_{\delta}+1)\big{)},$ (29)
and therefore
$r_{\rm{LT},\delta}(\tau_{\delta}+1)=q^{-1}\Big{(}(1+o(1))\frac{\log(1/\delta)}{J(y)}\Big{)}=J(y)^{-1/\beta}q^{-1}(\log 1/\delta)(1+o(1)).$ (30)
Recalling that $r_{\rm{LT},\delta}=q^{-1}(\log 1/\delta)$ by definition for light-tailed distibutions, we see that
$\tau_{\delta}+1=J(y)^{-1/\beta}(1+o(1)).$ (31)
Now, recall $y=(1-\gamma)y^{*}$, so that $J(y)=(1-\gamma)^{-\beta}J(y^{\ast})>1$ by (11), (9), as $J(y^{\ast})=1$. We conclude that $\tau_{\delta}\rightarrow-\gamma$ as $\delta\downarrow 0$.
Next we use that the constraint in ($\rm{CVaR}_{\delta}$) with $x=y/r_{\rm{LT},\delta}$ is equivalent to
${\bf E}\big{[}\phi(y,L)/r_{\rm{LT},\delta}\mid\phi(y,L)>r_{\rm{LT},\delta}(\tau_{\delta}+1)\big{]}\leq 1.$ (32)
We compute that
${\bf E}[\phi(y,L)/r_{\rm{LT},\delta}$ $\mid\phi(y,L)/r_{\rm{LT},\delta}>\tau_{\delta}+1]$
$\leq 1-\gamma/2+{\bf E}[(\phi(y,L)/r_{\rm{LT},\delta}-(1-\gamma/2))^{+}\mid\phi(y,L)/r_{\rm{LT},\delta}>\tau_{\delta}+1]$
$=1-\gamma/2+\frac{\int_{1-\gamma/2}^{\infty}{\bf P}(\phi(y,L)>r_{\rm{LT},\delta}u)\mathrm{d}u}{{\bf P}(\phi(y,L)>(1+\tau_{\delta})r_{\rm{LT},\delta})}.$
For fixed $y$, we can write ${\bf P}(\phi(y,L)>z)=\exp\{-\tilde{q}(z)\}$, with $\tilde{q}$ regularly varying of index $\beta>0$. Due to the Potter bounds (see e.g. Bingham et al. (1987), Theorem 1.5.6), for every $\varepsilon>0$ there exists a constant $z_{\varepsilon}$ such that, for $u\geq s\geq 1/2$, and $r\geq z_{\varepsilon}$
$\tilde{q}(ru)/\tilde{q}(rs)\geq(1-\varepsilon)(u/s)^{\beta-\varepsilon}.$ (33)
Using this bound, for sufficiently small $\varepsilon$ (w.r.t. $\gamma$ and $\beta$) and recalling that $\tau_{\delta}\rightarrow-\gamma$, it follows that
$\frac{\int_{1-\gamma/2}^{\infty}{\bf P}(\phi(y,L)>r_{\rm{LT},\delta}u)\mathrm{d}u}{{\bf P}(\phi(y,L)>(1+\tau_{\delta})r_{\rm{LT},\delta})}$ $\int_{1-\gamma/2}^{\infty}\exp\big{(}-\tilde{q}(r_{\rm{LT},\delta}u)+\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})\big{)}\mathrm{d}u$
$=\int_{1-\gamma/2}^{\infty}\exp\Big{(}-\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})\Big{(}\frac{\tilde{q}(r_{\rm{LT},\delta}u)}{\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})}-1\Big{)}\Big{)}\mathrm{d}u$
$\leq\int_{1-\gamma/2}^{\infty}\exp\Big{(}-\tilde{q}((1+\tau_{\delta})r_{\rm{LT},\delta})(1-\varepsilon)\Big{(}\big{(}\frac{u}{1+\tau_{\delta}}\big{)}^{\beta-\varepsilon}-1\Big{)}\Big{)}\mathrm{d}u$
$\rightarrow 0$
as $r_{\rm{LT},\delta}\to\infty$. We conclude that for every $\gamma>0$, the choice $y=y^{*}(1-\gamma)$ satisfies (32) for sufficiently small $\delta$, which implies that $\limsup_{\delta\downarrow 0}v_{{\bf CVaR},\delta}/v_{\delta}\geq 1-\gamma$. Theorem 2 now follows by letting $\gamma\downarrow 0$. $\square$
7.3. Proof of Theorem 3
Let $L^{(j)},j=1,...,k$ be $k$ i.i.d. samples of $L$. Combining (10) with Theorem 1 of Resnick and Tomkins (1973), we have that
$\frac{1}{q^{-1}(\log k)}\max_{j=1,...,k}b^{T}L^{(j)}\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\frac{1}{I(b)^{1/\beta}},\qquad k\to\infty.$ (34)
almost surely. We apply (34) to establish the following pointwise convergence for $y\geq 0$:
$g_{k}(y)$ $:=\frac{1}{q^{-1}(\log k)}\max_{j=1,...,k}\max_{i=1,...,d}y^{T}A^{(i)}L^{(j)}\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\max_{i=1,...,d}1/I(y^{T}A^{(i)})^{1/\beta}=:g(y).$ (35)
Observe that $g(y)$ is strictly increasing by (9). The statement follows now immediately from Proposition 1. $\square$
7.4. Proof of Theorem 4
We recall that $r_{\delta}=\bar{F}^{-1}(\delta)$. We first derive the pointwise limit of ${\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})$ as $\delta\downarrow 0$, using (12). Let $\varepsilon\in(0,1/\max_{i}y^{T}A^{(i)}e)$, with $e$ a vector of ones. As we assumed that the entries of all $A^{(i)}$ are non-negative, $|L|\leq\varepsilon r_{\delta}$ implies that $\phi(y,L)\leq r_{\delta}$. Consequently,
$\frac{{\bf P}(\phi(y,L)>r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ $=\frac{{\bf P}(\phi(y,L)>r_{\delta};|L|>\varepsilon r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ (36)
$={\bf P}\big{(}\phi(y,L)>r_{\delta}\,\big{|}\,|L|>\varepsilon r_{\delta}\big{)}\frac{{\bf P}(|L|>\varepsilon r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ (37)
$={\bf P}\Big{(}\phi\Big{(}y,\frac{L}{|L|}\Big{)}\frac{|L|}{r_{\delta}}>1\,\big{|}\,|L|>\varepsilon r_{\delta}\Big{)}\frac{{\bf P}(|L|>\varepsilon r_{\delta})}{{\bf P}(|L|>r_{\delta})}.$ (38)
Due to univariate regular variation, the ratio converges to $\varepsilon^{-\alpha}$. Due to the multivariate regular variation property (12), the conditional (on $|L|>\varepsilon r_{\delta}$) weak limit of $(L/|L|,|L|/r_{\delta})$ is $(\Theta,\varepsilon T)$, with $T$ a random variable with density $\alpha u^{-1-\alpha}$ independent of $\Theta$. Consequently,
$\frac{{\bf P}(\phi(y,L)>r_{\delta})}{{\bf P}(|L|>r_{\delta})}$ $\rightarrow\varepsilon^{-\alpha}{\bf P}(\phi(y,\Theta)\varepsilon T>1)$ (39)
$=\varepsilon^{-\alpha}\int_{1}^{\infty}\alpha u^{-1-\alpha}{\bf P}(\phi(y,\Theta)\varepsilon>1/u)\mathrm{d}u$ (40)
$=\int_{0}^{1/\varepsilon}\alpha v^{\alpha-1}{\bf P}(\phi(y,\Theta)>v)\mathrm{d}v$ (41)
$=\int_{0}^{\infty}\alpha v^{\alpha-1}{\bf P}(\phi(y,\Theta)>v)\mathrm{d}v$ (42)
$={\bf E}[\phi(y,\Theta)^{\alpha}].$ (43)
In the derivation above, the second equality follows from the change of variable $v=1/\varepsilon u$, the third equality comes from the fact that $\phi(y,\Theta)\leq\max_{i}y^{T}A^{(i)}e\leq 1/\varepsilon$, and the last equality from a representation for the $\alpha$th moment of a non-negative random variable.
To finish the proof, we define $k=1/\delta$, $g_{k}(y)={\bf P}(\phi(y,L)>r_{\delta})/{\bf P}(|L|>r_{\delta})$ and $g(y)={\bf E}[\phi(y,\Theta)^{\alpha}]$, and apply Proposition 1. $\square$
7.5. Proof of Theorem 5
The constraint in ($\rm{CVaR}_{\delta}$) with $x=y/r_{\delta}$ is equivalent to
${\bf E}\big{[}\phi(y,L)/r_{\delta}\mid\phi(y,L)>r_{\delta}(\tau_{\delta}+1)\big{]}\leq 1$ (44)
where $\tau_{\delta}$ is the optimizing value in (5). Using (12), it satisfies
$\delta={\bf P}(\phi(y/r_{\delta},L)-1>\tau_{\delta})$ $={\bf P}\big{(}\phi(y,L)>r_{\delta}(\tau_{\delta}+1)\big{)}$
$=(1+o(1)){\bf E}[\phi(y,\Theta)^{\alpha}]{\bf P}(|L|\geq r_{\delta}(\tau_{\delta}+1)).$
As a result, we obtain that $(\tau_{\delta}+1)r_{\delta}=\bar{F}^{-1}(\frac{(1+o(1))\delta}{{\bf E}[\phi(y,\Theta)^{\alpha}]})\to\infty$ as $\delta\downarrow 0$. We substitute this in the left-hand side in (44), and recall that $\bar{F}^{-1}(\delta)=r_{\delta}$ by definition. Then, as $\delta\downarrow 0$,
${\bf E}\big{[}\phi(y,L)/r_{\delta}\mid\phi(y,L)>r_{\delta}(\tau_{\delta}+1)\big{]}$ $=\frac{\bar{F}^{-1}\Big{(}\frac{(1+o(1))\delta}{{\bf E}[\phi(y,\Theta)^{\alpha}]}\Big{)}}{\bar{F}^{-1}(\delta)}\cdot{\bf E}\left[\frac{\phi(y,L)}{r_{\delta}(\tau_{\delta}+1)}\,\Big{|}\,\frac{\phi(y,L)}{r_{\delta}(\tau_{\delta}+1)}>1\right]$
$\rightarrow({\bf E}[\phi(y,\Theta)^{\alpha}])^{1/\alpha}\cdot\frac{\alpha}{\alpha-1}.$
The convergence of the first factor follows from regular variation. The convergence of the second factor holds because $\phi(y,L)/r_{\delta}(\tau_{\delta}+1)$, conditional upon exceeding 1, converges to a Pareto distributed random variable on $[1,\infty)$ with exponent $\alpha$. This convergence also holds in expectation; to show this we can apply a Potter bound for the ratio of regularly varying functions as we did in the proof of Theorem 2.
We now apply Proposition 1 with the values $k=1/\delta$, $g_{k}(y)={\bf E}[\phi(y,L)/r_{\delta}\mid\phi(y,L)>r_{\delta}(\tau_{\delta}+1)]$, and $g(y)=({\bf E}[\phi(y,\Theta)^{\alpha}])^{1/\alpha}\frac{\alpha}{\alpha-1}$. We conclude that the normalized version of ($\rm{CVaR}_{\delta}$) converges to the optimal value of
maximize $\mbox{maximize }c^{T}y\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }{\bf E}[\phi(y,\Theta)^{\alpha}]\leq(1-1/\alpha)^{\alpha}.$ (45)
Noting that ($\rm{HT}$) has a unique solution $y^{*}$, we conclude that the optimal solution to (45) equals $y^{*}(1-1/\alpha)$ by the homogeneity of $\phi$. Combining this with Theorem 4 completes the proof. $\square$
7.6. Proof of Theorem 6
Write $a_{k}:=\bar{F}^{-1}(1/k)$ for the normalizing constants in (${\rm{SCP}\textnormal{-}\rm{HT}}_{k}$). To show convergence in distribution of the optimal value $V_{SCP}(k)$ to $V^{\ast}$, it suffices to show that the non-feasible set spanned by the constraints $\phi(y,L^{(i)}/a_{k})\leq 1$, $i=1,\ldots,k$, converges weakly to the non-feasible set spanned by the constraints $\phi(y,z)\leq 1$, $z\in N$, where $N$ is the Poisson point process defined above (14). To do so, we rely on the theory of random sets, as developed in Molchanov (2017).
To translate our setting to that of Section 4.2 in Molchanov (2017), define the multi-function $M(L):=\{y:\phi(y,L)\geq 1\}$. By the homogeneity of $\phi$, the closure of the set of non-feasible solutions for (${\rm{SCP}\textnormal{-}\rm{HT}}_{k}$) can be written as
$U_{k}=\bigcup_{i=1}^{k}\Big{\{}y:\phi\Big{(}y,\frac{L^{(i)}}{a_{k}}\Big{)}\geq 1\Big{\}}=\frac{1}{a_{k}}\bigcup_{i=1}^{k}M(L^{(i)}),$ (46)
where we use the notation $bS=\{x:bx\in S\}$. Applying Theorem 4.2.9 and 4.2.10 of Molchanov (2017), we obtain that $U_{k}$ converges in distribution to $U=\{y:\phi(y,z)\geq 1\mbox{ for some }z\in N\}$. We define for $t>0$ the half-space $K_{t}=\{y:c^{T}y>t\}$. Therefore, we see that for $t>0$,
${\bf P}(V_{SCP}(k)\leq t)$ $={\bf P}\big{(}\forall y\in K_{t}\exists i\in\{1,\ldots,k\}:\phi\big{(}y,L^{(i)}/a_{k}\big{)}>1\big{)}$ (47)
$={\bf P}\big{(}\forall y\in K_{t}\exists i\in\{1,\ldots,k\}:\phi\big{(}y,L^{(i)}/a_{k}\big{)}\geq 1\big{)}$ (48)
$={\bf P}(U_{k}\cap K_{t}\neq\emptyset)\rightarrow{\bf P}(U\cap K_{t}\neq\emptyset)={\bf P}(V^{*}\leq t).$ (49)
This completes the proof of Theorem 6. $\square$